Please typed I can copy and past, not hand writing! Thanks

Question 3 (4 marks) Part a) Given the following information X-500, σ=12, n=50 i) Determine the 95% confidence interval estimate of population mean. ii) Determine the 99% o confidence interval estimate of population mean Part b) A statistics practitioner calculated the mean and standard deviation from a sample of 51. They are X-120 and s-15. (i) Estimate the population mean with 95% confidence level (ii) Estimate the population mean with 99% confidence level.
Answer #1

Answer:

Given that:

Part (a) (i)   X = 500 , 12 , n= 50

We know that

The confident level is 95%.

Here ar{X} - E < mu < ar{X} + E

We consider

Tt

e=95% ,   a = 0.05

Using critical value distribution   Z_{c} =1.96

Tt

  12 E = 1.96 ×

= 1.96 x 1.69705

   =  3.3262

  X-E = 500-3.3262 496.6738

  ar{X+} E= 500+3.3262=503.3262.

The confident level is (496.6738 ,503.3262)

(ii) To find the 99% of confidence level

We know that  a-0.01Zo = 2.575

  

Tt

E=2.575 imes rac{12}{sqrt{50}}

E=2.575 imes 1.6970

  E=4.3699

X-E = 500-4.3699 495.6301
  X + E 500 + 4.3699 = 504.3699

The 99% of confidence level is (495.6301 ,504.3699)

Part (b) (i) We know that X 120, S 15, n-51

95% of confident interval to estimated

Here ar{X} - E < mu < ar{X} + E

E=t_{c} imes rac{S}{sqrt{n}}

Degree of freedom = 51-1=50 , alpha =0.05

Critical value for t-distribution t_{c}=2.00g

   E=t_{c} imes rac{S}{sqrt{n}}

  E=2.00 imes rac{15}{sqrt{51}}

=2.00 imes 2.1004

E=4.2197
ar{X}-E=120.4.2197=115.7803

  ar{X}+E=120+4.2197=124.2197

95% of confident level is (115.7803,124.2197)

(ii) 99% of confident level

alpha =0.01 , t_{c}= 2.678

     E=t_{c} imes rac{S}{sqrt{n}}

E=2.678 imes rac{15}{sqrt{51}}

  =2.678 imes 2.1004

  E=5.6249

ar{X}-E=120-5.6249=114.3751   

X + E 120 + 5.6249 125.6249

The 99% confident interval is (114.3751,125.6249)